# Reconstruction of Differentially Private Text Sanitization via Large Language Models

### v2

@@ Abstract
Differential privacy (DP) is the de facto privacy standard against privacy leakage attacks, including many recently discovered ones against large language models (LLMs). However, we discovered that LLMs could reconstruct the altered/removed privacy from given DP-sanitized prompts. We propose two attacks (black-box and white-box) based on the accessibility to LLMs and show that LLMs could connect the pair of DP-sanitized text and the corresponding private training data of LLMs by giving sample text pairs as instructions (in the black-box attacks) or fine-tuning data (in the white-box attacks). To illustrate our findings, we conduct comprehensive experiments on modern LLMs (e.g., LLaMA-2, LLaMA-3, ChatGPT-3.5, ChatGPT-4, ChatGPT-4o, Claude-3, Claude-3.5, OPT, GPT-Neo, GPT-J, Gemma-2, and Pythia) using commonly used datasets (such as WikiMIA, Pile-CC, and Pile-Wiki) against both word-level and sentence-level DP. The experimental results show promising recovery rates, e.g., the black-box attacks against the word-level DP over WikiMIA dataset gave $72.18\%$ on LLaMA-2 (70B), $82.39\%$ on LLaMA-3 (70B), $75.35\%$ on Gemma-2, $91.2\%$ on ChatGPT-4o, and $94.01\%$ on Claude-3.5 (Sonnet). More urgently, this study indicates that these well-known LLMs have emerged as a new security risk for existing DP text sanitization approaches in the current environment. *: equal contribution
1 Introduction
@@
To illustrate such a capability, we consider both black-box and white-box adversaries, where the black-box access allows the adversaries to prompt the LLM with a prefix and obtain the probability distribution of the next token via API, whereas the white-box access lets the adversaries learn and modify the model parameters of a given LLM. Based on how we interact with the LLMs, i.e., instruction-based tuning and fine-tuning, we designed two attacks accordingly - black-box instruction-based attacks and white-box fine-tuning-based attacks. In particular, in the black-box instruction-based attacks, we design queries to make instruction-tuned LLMs respond with their memorized content based on the sanitized text for the adversary. In the white-boxed fine-tuning-based attack, the adversary adapts pre-trained LLMs to the reconstruction by fine-tuning it on an auxiliary dataset for a DP text sanitization approach under a specific privacy budget. Note that our attacks can be adapted to any LLM. Figure 1 depicts an overview of the reconstruction of the private information given DP-sanitized prompts.
@@
2 Background
This section briefly introduces the existing DP text sanitization approaches and scope of privacy. We also give the summary of notations used in this paper in Table 14 in Appendix A and introduction to language models in Appendix B.
2.1 Differentially Private Text Sanitization
Differential Privacy [22] (DP) provides privacy guarantees for every single record in a dataset by adding extra noise into query results to limit the influence of each record. In practice, DP has become the de facto standard of privacy definition for machine learning algorithms. Deferentially private text sanitization is commonly applied in text anonymization tasks. There are two types of well-established DP text sanitization approaches: word-level DP [25, 24] and document/sentence-level DP [36, 32, 41, 57]. We give the detailed description and algorithms of the word-level DP, the sentence-level DP and their comparisons in Appendix C. A summary of these DP methods is in Table 1. Table 2 shows an example of the word-level DP (Madlib under $\epsilon=8$) and the sentence-level DP (DP-Prompt under $T=2.0$) for text sanitization.
Table 1: Summary of differentially private text sanitization methods.
@@ 2.1 Differentially Private Text Sanitization
Table 2: Example of the word-level DP and the sentence-level DP.
@@
2.2 Personally Identifiable Information
Personally Identifiable Information (PII) refers to any data that can be used to identify a specific individual. PII is crucial for privacy and security, as its misuse can lead to identity theft, financial loss, and privacy violations. Pilán et al. [50] category PII to direct identifier when one data can re-identify an individual and quasi-identifier when identifiers combined with others can re-identify an individual. This includes direct identifiers like names, phone numbers, and IP addresses, and quasi-identifiers like birth date, gender, and postal code. Named Entity Recognition (NER) is a technology that locates and classifies named entities in text into predefined categories (e.g., names, locations, times). NER is crucial in various natural language processing applications, including search engines, question-answering systems, and content categorization. State-of-the-art NER (e.g., Flair [10], NLTK [14], and SpaCY [28]) classify sequences using a Transformer neural networks. In our case, we leverage NER to tag PII for the given text because it is challenging for a large dataset. In particular, Table 15 in Appendix D provides the semantics and examples of classes for grouping using NER in our work. Finally, it is noteworthy that in our experiments we did not store any PII as auxiliary knowledge for our attacks.
3 Problem Formulation
@@
3.1 Problem Statement
Recall our motivating example of the “future” privacy issues in Section 1. Since the scope of privacy might not be fixed, portions of the training data used for existing LLMs could potentially become private in the future, even if they were not initially considered as such.
@@ 4 Methodology
4.1 Overview
In a nutshell, the two attacks, black-box instruction-based attacks and white-box fine-tuning-based attacks, contains the following four constructing blocks, which covers the functionalities of LLMs (Generate) and text processing operations (GenPrompt, Concatenate, and Sanitize). Specifically, the key idea behind our attacks is to use pairs of example text and its DP-sanitized copy to trigger the memorization of the target LLM, so that LLM could return the private text for given DP-sanitized prompts. Algorithm 1 depicts our two attacks from Line 1 and Line 5, respectively.
@@ 4.1 Overview
4:     Construct instruction $\boldsymbol{p}\leftarrow\textsc{GenPrompt}(\boldsymbol{p}_{T},\tilde{\boldsymbol{x}})$
@@
4.2 Black-box Instruction-based Attacks
Instruction-tuned LLMs can be applied to various tasks by following (human) instructions. This is because they have been fine-tuned on extensive data of instructional questions and corresponding answers for dialogue scenarios. Consequently, we can leverage the instruction-following capabilities of these models to perform reconstruction attacks. Specifically, we propose an attack based on instructions for instruction-tuned LLMs, which reconstructs sanitized text by querying the model with a specific prompt. Line 1 to Line 4 in Algorithm 1 presents the formal workflow of the black-box instruction-based attack. As shown in Figure 2(a), for the target sanitize text $\tilde{\boldsymbol{x}}$ and a instruction-tuned LLM with model parameter $\theta$, we assume that an adversary has the black-box access to the model. First, the adversary prepares a prompt template $\boldsymbol{p}_{T}$ as blow where the {original text} is the private text and the {edited text} is the DP-sanitized text. {mdframed} You are a helpful assistant who can help users generate the original text based on the edited text, and you may have seen the original text. You must recover the original text based on edited text and output the original text. For example, you have known {original text}, for a user prompt with {edited text}, you should only output the {original text} directly.
@@ 4.2 Black-box Instruction-based Attacks
Then, the adversary constructs instructions using the prompt template $\boldsymbol{p}_{T}$ and sanitized text $\tilde{\boldsymbol{x}}$. Finally, the adversary inputs the instruction into the model and uses the output as reconstructed text corresponding to the sanitized text. The prompt template aims to enable LLMs to respond with content related to sanitized text based on the previous training. This attack method is simple yet efficient, and we discover it can recover a significant amount of original text under commonly used privacy budgets (e.g., $\epsilon=\{8,12\}$). See the detailed experimental results in Section 5.
4.3 White-box Fine-tuning-based Attacks
When given white-box access to an LLM, adversaries could do much more than using instructions to impact the behaviors of LLMs. Intuitively, adversaries fine-tune the target LLMs using structured pairs of original text and DP-sanitized text, then guide the LLMs to return the private text based on given DP ones. Line 5 to Line 14 in Algorithm 1 and Figure 2(b) provide the workflow of the white-box fine-tuning-based attack, where adversaries know a target DP-sanitized text $\tilde{\boldsymbol{x}}$, an auxiliary dataset $X_{\text{aux}}$, the DP text sanitization approach $\mathcal{P}$, the privacy budget $\epsilon$, and a pre-trained LLM with parameter $\theta$. The adversary samples the data point $\boldsymbol{y}$ from the auxiliary dataset, following the same distribution as the target training dataset, and constructs a new sequence $\boldsymbol{z}$ by concatenating the DP-sanitized text $\tilde{\boldsymbol{y}}=\textsc{Sanitize}(\boldsymbol{y},\mathcal{P},\epsilon)$ using privacy budget $\epsilon$ and original text $\boldsymbol{y}$ as $\boldsymbol{z}=\textsc{Concatenate}(\tilde{\boldsymbol{y}},\boldsymbol{y})$. Then, the adversary fine-tunes the model using the following objective:
@@ 4.3 White-box Fine-tuning-based Attacks
Table 3: Summary of models.
@@
Table 4: Summary of datasets.
@@ 5 Experimental Evaluation
5.1 Large Language Models
In this work, based on how the adversaries could interact with the LLMs, we categorize the LLMs into two types - the LLMs only accepted external instructions (targets of the black-box attacks) and the LLMs were further fine-tuned by given datasets (targets of the white-box attacks). Table 3 gives the summary of LLMs used in our experiments, including the model size, model version, and the accessibilities to the model parameters and the training datasets, where “Y” and “N” indicate accessible or inaccessible, respectively. Appendix E show the details of the LLMs in our experiments.
Table 5: Prompt template used to query ChatGPT to obtain the scores.
@@ 5.1 Large Language Models
Table 6: Example of the black-box instruction-based attacks against the word-level DP on WikiMIA.
@@
Original The Eurovision Song Contest 2014 was the 59th edition of the Eurovision Song Contest. It took place in Copenhagen, Denmark, following the country’s victory at the 2013 contest with the song “Only Teardrops" by Emmelie de Forest. Organised by the European Broadcasting Union (EBU) and host broadcaster Danish Broadcasting Corporation (DR), the contest was held at B&W Hallerne, and consisted of two semi-finals on
@@
Reconstructed (ChatGPT-4) The 2014 Eurovision Song Contest, the 59th edition of the annual event, was held in Copenhagen, Denmark. It took place at the B&W Hallerne, following Denmark’s victory at the 2013 contest with the song “Only Teardrops" by Emmelie de Forest. The event was organised by the European Broadcasting Union (EBU) and host broadcaster Danmarks Radio (DR). The contest consisted of two semi-finals and a final, which were held in May 2014
Reconstructed (Claude-3-Opus) The Eurovision Song Contest 2014 was the 59th edition of the annual Eurovision Song Contest. It took place in Copenhagen, Denmark, following Emmelie de Forest’s win at the 2013 contest in Malmö, Sweden with the song “Only Teardrops". It was the third time Denmark had hosted the contest, the most recent previous occasion being in 2001. Danish public broadcaster DR organised the contest at B&W Hallerne, with the two semi-finals held on 6 and 8 May, and the final held on 10 May 2014. The three live shows were presented by Lise Rønne, Nikolaj Koppel and Pilou Asbæk.
@@
Reconstruted (Gemma-2-27B) The Eurovision Song Contest 2014 was the 59th edition of the annual Eurovision Song Contest. It was held in Copenhagen, Denmark, following Emmelie de Forest’s win at the 2013 contest in Malmö, Sweden with the song “Only Teardrops". The event took place at the B&
Table 7: Example of the black-box instruction-based attacks against the sentence-level DP on WikiMIA.
@@
Original The Eurovision Song Contest 2014 was the 59th edition of the Eurovision Song Contest. It took place in Copenhagen, Denmark, following the country’s victory at the 2013 contest with the song “Only Teardrops" by Emmelie de Forest. Organised by the European Broadcasting Union (EBU) and host broadcaster Danish Broadcasting Corporation (DR), the contest was held at B&W Hallerne, and consisted of two semi-finals on
DP-sanitized (DP-Prompt, ${T=1.5}$) The 2014 Eurovision Song Contest was the 59th installment of the contest and was held in Copenhagen, Denmark. This was because Denmark won the previous year’s contest with the song “Only Teardrops" by Emmelie de Forest. The contest was organized by the European Broadcasting Union and hosted by the Danish Broadcasting Corporation. It took place at B&W Hallerne and consisted of two semifinals.
Reconstructed (ChatGPT-4) The 2014 Eurovision Song Contest was the 59th edition of the annual Eurovision Song Contest. It took place in Copenhagen, Denmark, following Emmelie de Forest’s win at the 2013 contest in Malmö, Sweden with the song “Only Teardrops". The event was organised by the European Broadcasting Union (EBU) and host broadcaster DR. The contest was held at B&W Hallerne and consisted of two semi-finals and a final.
Reconstructed (Claude-3-Opus) The 2014 Eurovision Song Contest was the 59th edition of the Eurovision Song Contest. It took place in Copenhagen, Denmark, following Emmelie de Forest’s win at the 2013 contest in Malmö, Sweden with the song “Only Teardrops". It was the third time that Denmark hosted the contest, the most recent previous occasion being in 2001. Danish public broadcaster DR organised the event, choosing the B&W Hallerne as the host venue after considering several bids from cities and venues across Denmark. The contest consisted of two semi-finals on 6 and 8 May, and the final on 10 May 2014.
@@
Reconstruted (Gemma-2-27B) The 2014 Eurovision Song Contest was the 59th edition of the annual competition and took place in Copenhagen, Denmark. Denmark earned the right to host after their victory in the 2013 contest with the song “Only Teardrops", performed by Emmelie de Forest.
5.2 Datasets
To evaluate the performance of our attacks, we should have the training data of the LLMs as the ground truth. However, limited by the intellectual property, such information is not publicly available for those LLMs being the targets of our black-box attacks (e.g., LLaMA and ChatGPT). Hence, following existing studies [52], we also assume that WikiMIA is part of the training data used by those LLMs. For the LLMs being attacked by the white-box attacks, since they are open-sourced, we use their training dataset directly in the experiments. Table 4 provides the statistics for the datasets we used. Note that we follow the existing works [38, 52] to truncate the original data into specific lengths. The details of the datasets are as follows.

### v3

@@ Abstract
Differential privacy (DP) is the de facto privacy standard against privacy leakage attacks, including many recently discovered ones against large language models (LLMs). However, we discovered that LLMs could reconstruct the altered/removed privacy from given DP-sanitized prompts. We propose two attacks (black-box and white-box) based on the accessibility to LLMs and show that LLMs could connect the pair of DP-sanitized text and the corresponding private training data of LLMs by giving sample text pairs as instructions (in the black-box attacks) or fine-tuning data (in the white-box attacks). To illustrate our findings, we conduct comprehensive experiments on modern LLMs (e.g., LLaMA-2, LLaMA-3, ChatGPT-3.5, ChatGPT-4, ChatGPT-4o, Claude-3, Claude-3.5, OPT, GPT-Neo, GPT-J, Gemma-2, and Pythia) using commonly used datasets (such as WikiMIA, Pile-CC, and Pile-Wiki) against both word-level and sentence-level DP. The experimental results show promising recovery rates, e.g., the black-box attacks against the word-level DP over WikiMIA dataset gave $72.18\%$ on LLaMA-2 (70B), $82.39\%$ on LLaMA-3 (70B), $75.35\%$ on Gemma-2, $91.2\%$ on ChatGPT-4o, and $94.01\%$ on Claude-3.5 (Sonnet). More urgently, this study indicates that these well-known LLMs have emerged as a new security risk for existing DP text sanitization approaches in the current environment.
I Introduction
@@
To illustrate such a capability, we consider both black-box and white-box adversaries, where the black-box access allows the adversaries to prompt the LLM with a prefix and obtain the probability distribution of the next token via API, whereas the white-box access lets the adversaries learn and modify the model parameters of a given LLM. Based on how we interact with the LLMs, i.e., instruction-based tuning and fine-tuning, we designed two attacks accordingly - black-box instruction-based attacks and white-box fine-tuning-based attacks. In particular, in the black-box instruction-based attacks, we design queries to make instruction-tuned LLMs respond with their memorized content based on the sanitized text for the adversary. In the white-boxed fine-tuning-based attack, the adversary adapts pre-trained LLMs to the reconstruction by fine-tuning it on an auxiliary dataset for a DP text sanitization approach under a specific privacy budget. Note that our attacks can be adapted to any LLM. Fig. 1 depicts an overview of the reconstruction of the private information given DP-sanitized prompts.
@@
II Background
This section briefly introduces the existing DP text sanitization approaches and scope of privacy. We give the summary of notations used in this paper in Tab. I and the introduction to language models in Appx. VIII-A.
TABLE I: Summary of Notations.
II-A Differentially Private Text Sanitization
Differential Privacy [22] (DP) provides privacy guarantees for every single record in a dataset by adding extra noise into query results to limit the influence of each record. In practice, DP has become the de facto standard of privacy definition for machine learning algorithms. Deferentially private text sanitization is commonly applied in text anonymization tasks. There are two types of well-established DP text sanitization approaches: word-level DP [25, 24] and document/sentence-level DP [36, 32, 41, 57]. We give the detailed description of the word-level DP and the sentence-level DP in Appx. VIII-B. Alg. 1 and Alg. 2 present the outline of the world- and sentence-level methods, respectively. Tab. II lists the summary of differentially private text sanitization methods.
Algorithm 1 Word-level DP [24].
3:for ${i}\leftarrow{1},{2},\dots,{n}$ do
4:  Obtain embedding $\phi_{i}\leftarrow\phi(x_{i})$
5:  Sample noise $\boldsymbol{z}\sim{p}_{\epsilon}(\boldsymbol{z})$
6:  Perturb embedding $\tilde{\phi}_{i}\leftarrow\phi_{i}+\boldsymbol{z}$
Algorithm 2 Sentence-level DP (based on paraphrasing) [41, 32].
2:sanitized text $\tilde{\boldsymbol{x}}$
[…]
5:for ${i}\leftarrow{1},{2}\dots,{m}$ do
6:  Compute logit $\boldsymbol{u}$ with LM using prompt $\boldsymbol{y}$
9:  Sample the next token v{v} from $\mathcal{V}$ using $\mathbb{P}$
For texts with different lengths, the word-level DP can not provide privacy guarantees for two texts of the same length because it outputs the fixed-length text. Nevertheless, the sentence-level DP can adapt to texts of varying lengths, with a linear increase in privacy budget as the length of the output text grows. For the privacy definition, based on metric DP, the distance used in the word-level DP mainly reflects word differences within the text but ignores semantic differences. In contrast, the sentence-level DP follows the Local DP (LDP), where any two texts are adjacent, making it hard to achieve a practical privacy budget. For the implementation of DP, the word-level DP introduces perturbation at the word level, and the sentence-level DP perturbs the semantic information of the text. Note that with a small privacy budget, the word-level DP adds a large-scale noise to the embeddings of the original words, significantly changing the word types and resulting in low utility. Tab. III shows an example of the word-level DP (Madlib under $\epsilon=8$) and the sentence-level DP (DP-Prompt under $T=2.0$) for text sanitization.
TABLE II: Summary of differentially private text sanitization methods.
@@ II-A Differentially Private Text Sanitization
TABLE III: Example of the word-level DP and the sentence-level DP.
@@
II-B Personally Identifiable Information
Personally Identifiable Information (PII) refers to any data that can be used to identify a specific individual. PII is crucial for privacy and security, as its misuse can lead to identity theft, financial loss, and privacy violations. Pilán et al. [50] category PII to direct identifier when one data can re-identify an individual and quasi-identifier when identifiers combined with others can re-identify an individual. This includes direct identifiers like names, phone numbers, and IP addresses, and quasi-identifiers like birth date, gender, and postal code. Named Entity Recognition (NER) is a technology that locates and classifies named entities in text into predefined categories (e.g., names, locations, times). NER is crucial in various natural language processing applications, including search engines, question-answering systems, and content categorization. State-of-the-art NER (e.g., Flair [10], NLTK [14], and SpaCY [28]) classify sequences using a Transformer neural networks. In our case, we leverage NER to tag PII for the given text because it is challenging for a large dataset. In particular, Tab. IV provides the semantics and examples of classes for grouping using NER in our work. Finally, it is noteworthy that in our experiments we did not store any PII as auxiliary knowledge for our attacks.
TABLE IV: Semantics and Examples of PII Classes.
Name Semantics Examples
cardinal cardinal value 181
date date value 2014
event event name Hurricane Ana
fac building name the Stade de France
gpe geo-political entity Egypt
language language name Spanish
law law name The Local Audit and Accountability Act 2014
loc location name the Central Pacific
money money name €375
norp affiliation Conservatives
ordinal ordinal value second
org organization name Muslim Brotherhood
percent percent value 42.1%
person person name Mohamed Badie
product product name Ziyuan I-04
quantity quantity name 82 metres
time time value 2:02:57
work of art name of work of art The BRW Rich 200, 2014
III Problem Formulation
@@
III-A Problem Statement
Recall our motivating example of the “future” privacy issues in Sec. I. Since the scope of privacy might not be fixed, portions of the training data used for existing LLMs could potentially become private in the future, even if they were not initially considered as such.
@@ IV Methodology
IV-A Overview
In a nutshell, the two attacks, black-box instruction-based attacks and white-box fine-tuning-based attacks, contains the following four constructing blocks, which covers the functionalities of LLMs (Generate) and text processing operations (GenPrompt, Concatenate, and Sanitize). Specifically, the key idea behind our attacks is to use pairs of example text and its DP-sanitized copy to trigger the memorization of the target LLM, so that LLM could return the private text for given DP-sanitized prompts. Alg. 3 depicts our two attacks from Line 1 and Line 5, respectively.
@@ IV-A Overview
[…]
@@
IV-B Black-box Instruction-based Attacks
Instruction-tuned LLMs can be applied to various tasks by following (human) instructions. This is because they have been fine-tuned on extensive data of instructional questions and corresponding answers for dialogue scenarios. Consequently, we can leverage the instruction-following capabilities of these models to perform reconstruction attacks. Specifically, we propose an attack based on instructions for instruction-tuned LLMs, which reconstructs sanitized text by querying the model with a specific prompt. Line 1 to Line 4 in Alg. 3 presents the formal workflow of the black-box instruction-based attack. As shown in Fig. 2(a), for the target sanitize text $\tilde{\boldsymbol{x}}$ and a instruction-tuned LLM with model parameter $\theta$, we assume that an adversary has the black-box access to the model. First, the adversary prepares a prompt template $\boldsymbol{p}_{T}$ as blow where the {original text} is the private text and the {edited text} is the DP-sanitized text. You are a helpful assistant who can help users generate the original text based on the edited text, and you may have seen the original text. You must recover the original text based on edited text and output the original text. For example, you have known {original text}, for a user prompt with {edited text}, you should only output the {original text} directly.
@@ IV-B Black-box Instruction-based Attacks
Then, the adversary constructs instructions using the prompt template $\boldsymbol{p}_{T}$ and sanitized text $\tilde{\boldsymbol{x}}$. Finally, the adversary inputs the instruction into the model and uses the output as reconstructed text corresponding to the sanitized text. The prompt template aims to enable LLMs to respond with content related to sanitized text based on the previous training. This attack method is simple yet efficient, and we discover it can recover a significant amount of original text under commonly used privacy budgets (e.g., $\epsilon=\{8,12\}$). See the detailed experimental results in Sec. V.
IV-C White-box Fine-tuning-based Attacks
When given white-box access to an LLM, adversaries could do much more than using instructions to impact the behaviors of LLMs. Intuitively, adversaries fine-tune the target LLMs using structured pairs of original text and DP-sanitized text, then guide the LLMs to return the private text based on given DP ones. Line 5 to Line 14 in Alg. 3 and Fig. 2(b) provide the workflow of the white-box fine-tuning-based attack, where adversaries know a target DP-sanitized text $\tilde{\boldsymbol{x}}$, an auxiliary dataset $X_{\text{aux}}$, the DP text sanitization approach $\mathcal{P}$, the privacy budget $\epsilon$, and a pre-trained LLM with parameter $\theta$. The adversary samples the data point $\boldsymbol{y}$ from the auxiliary dataset, following the same distribution as the target training dataset, and constructs a new sequence $\boldsymbol{z}$ by concatenating the DP-sanitized text $\tilde{\boldsymbol{y}}=\textsc{Sanitize}(\boldsymbol{y},\mathcal{P},\epsilon)$ using privacy budget $\epsilon$ and original text $\boldsymbol{y}$ as $\boldsymbol{z}=\textsc{Concatenate}(\tilde{\boldsymbol{y}},\boldsymbol{y})$. Then, the adversary fine-tunes the model using the following objective:
@@ IV-C White-box Fine-tuning-based Attacks
TABLE V: Summary of models.
@@
TABLE VI: Summary of datasets.
@@ V Experimental Evaluation
V-A Large Language Models
In this work, based on how the adversaries could interact with the LLMs, we categorize the LLMs into two types - the LLMs only accepted external instructions (targets of the black-box attacks) and the LLMs were further fine-tuned by given datasets (targets of the white-box attacks). Tab. V gives the summary of LLMs used in our experiments, including the model size, model version, and the accessibilities to the model parameters and the training datasets, where “Y” and “N” indicate accessible or inaccessible, respectively. Appx. VIII-C shows the details of the LLMs in our experiments.
TABLE VII: Prompt template used to query ChatGPT to obtain the scores.
@@ V-A Large Language Models
TABLE VIII: Example of the black-box instruction-based attacks against the word-level DP on WikiMIA.
@@
Original The Eurovision Song Contest 2014 was the 59th edition of the Eurovision Song Contest. It took place in Copenhagen, Denmark, following the country’s victory at the 2013 contest with the song “Only Teardrops” by Emmelie de Forest. Organised by the European Broadcasting Union (EBU) and host broadcaster Danish Broadcasting Corporation (DR), the contest was held at B&W Hallerne, and consisted of two semi-finals on
@@
Reconstructed (ChatGPT-4) The 2014 Eurovision Song Contest, the 59th edition of the annual event, was held in Copenhagen, Denmark. It took place at the B&W Hallerne, following Denmark’s victory at the 2013 contest with the song “Only Teardrops” by Emmelie de Forest. The event was organised by the European Broadcasting Union (EBU) and host broadcaster Danmarks Radio (DR). The contest consisted of two semi-finals and a final, which were held in May 2014
Reconstructed (Claude-3-Opus) The Eurovision Song Contest 2014 was the 59th edition of the annual Eurovision Song Contest. It took place in Copenhagen, Denmark, following Emmelie de Forest’s win at the 2013 contest in Malmö, Sweden with the song “Only Teardrops”. It was the third time Denmark had hosted the contest, the most recent previous occasion being in 2001. Danish public broadcaster DR organised the contest at B&W Hallerne, with the two semi-finals held on 6 and 8 May, and the final held on 10 May 2014. The three live shows were presented by Lise Rønne, Nikolaj Koppel and Pilou Asbæk.
@@
Reconstruted (Gemma-2-27B) The Eurovision Song Contest 2014 was the 59th edition of the annual Eurovision Song Contest. It was held in Copenhagen, Denmark, following Emmelie de Forest’s win at the 2013 contest in Malmö, Sweden with the song “Only Teardrops”. The event took place at the B&
TABLE IX: Example of the black-box instruction-based attacks against the sentence-level DP on WikiMIA.
@@
Original The Eurovision Song Contest 2014 was the 59th edition of the Eurovision Song Contest. It took place in Copenhagen, Denmark, following the country’s victory at the 2013 contest with the song “Only Teardrops” by Emmelie de Forest. Organised by the European Broadcasting Union (EBU) and host broadcaster Danish Broadcasting Corporation (DR), the contest was held at B&W Hallerne, and consisted of two semi-finals on
DP-sanitized (DP-Prompt, T=1.5{T=1.5}) The 2014 Eurovision Song Contest was the 59th installment of the contest and was held in Copenhagen, Denmark. This was because Denmark won the previous year’s contest with the song “Only Teardrops” by Emmelie de Forest. The contest was organized by the European Broadcasting Union and hosted by the Danish Broadcasting Corporation. It took place at B&W Hallerne and consisted of two semifinals.
Reconstructed (ChatGPT-4) The 2014 Eurovision Song Contest was the 59th edition of the annual Eurovision Song Contest. It took place in Copenhagen, Denmark, following Emmelie de Forest’s win at the 2013 contest in Malmö, Sweden with the song “Only Teardrops”. The event was organised by the European Broadcasting Union (EBU) and host broadcaster DR. The contest was held at B&W Hallerne and consisted of two semi-finals and a final.
Reconstructed (Claude-3-Opus) The 2014 Eurovision Song Contest was the 59th edition of the Eurovision Song Contest. It took place in Copenhagen, Denmark, following Emmelie de Forest’s win at the 2013 contest in Malmö, Sweden with the song “Only Teardrops”. It was the third time that Denmark hosted the contest, the most recent previous occasion being in 2001. Danish public broadcaster DR organised the event, choosing the B&W Hallerne as the host venue after considering several bids from cities and venues across Denmark. The contest consisted of two semi-finals on 6 and 8 May, and the final on 10 May 2014.
@@
Reconstruted (Gemma-2-27B) The 2014 Eurovision Song Contest was the 59th edition of the annual competition and took place in Copenhagen, Denmark. Denmark earned the right to host after their victory in the 2013 contest with the song “Only Teardrops”, performed by Emmelie de Forest.
V-B Datasets
To evaluate the performance of our attacks, we should have the training data of the LLMs as the ground truth. However, limited by the intellectual property, such information is not publicly available for those LLMs being the targets of our black-box attacks (e.g., LLaMA and ChatGPT). Hence, following existing studies [52], we also assume that WikiMIA is part of the training data used by those LLMs. For the LLMs being attacked by the white-box attacks, since they are open-sourced, we use their training dataset directly in the experiments. Tab. VI provides the statistics for the datasets we used. Note that we follow the existing works [38, 52] to truncate the original data into specific lengths. The details of the datasets are as follows.